# Streaming Sequence-to-Sequence Learning with Delayed Streams Modeling

### v1

@@ Abstract
We introduce Delayed Streams Modeling (DSM), a flexible formulation for streaming, multimodal sequence-to-sequence learning. Sequence-to-sequence generation is often cast in an offline manner, where the model consumes the complete input sequence before generating the first output timestep. Alternatively, streaming sequence-to-sequence rely on learning a policy for choosing when to advance on the input stream, or write to the output stream. DSM instead models already time-aligned streams with a decoder-only language model. By moving the alignment to a pre-processing step, and introducing appropriate delays between streams, DSM provides streaming inference of arbitrary output sequences, from any input combination, making it applicable to many sequence-to-sequence problems. In particular, given text and audio streams, automatic speech recognition (ASR) corresponds to the text stream being delayed, while the opposite gives a text-to-speech (TTS) model. We perform extensive experiments for these two major sequence-to-sequence tasks, showing that DSM provides state-of-the-art performance and latency while supporting arbitrary long sequences, being even competitive with offline baselines. Code, samples and demos are available at github.com/kyutai-labs/delayed-streams-modeling.
@@ 4.2 Training protocol
Finetuning (DSM-ASR). We then finetune the model on a collection of public datasets with ground-truth transcripts, described in Appendix A.1 and totaling 24k hours. This training stage lasts for 100k updates with batches of 128 examples, using 16 H100s. We augment training by using the codebook dropout (Défossez et al., 2024). We then adapt the model to long-form inputs, which most public datasets lack, by constructing a long-form mixture described in Appendix A.2. We run this stage for DSM-ASR for 25k updates with batch size 32, using 16 H100s.
@@ 4.3 Automatic Speech Recognition
Baselines. We benchmark DSM-ASR against leading models of the OpenASR Leaderboard, including Whisper (Radford et al., 2023), Canary-Flash (Zelasko et al., 2025), Phi-4 Multimodal Instruct (Abouelenin et al., 2025), Parakeet (Xu et al., 2023), Voxtral Mini (Liu et al., 2025) and as well as the closed-source ElevenLabs API 111elevenlabs.io.. Notably, all these models perform non-streaming ASR, as they require access to the full input sequence. We thus also include a streaming variant of Whisper (Macháček et al., 2023), with a delay of 2.5s and SeamlessStreaming (Barrault et al., 2023). For long-form transcription, we add the Distil-Whisper (Gandhi et al., 2023) variant.
Transcription quality. We report micro-averaged Word Error Rate (WER), which avoids overemphasizing short sequences, and is the standard computation used in the OpenASR Leaderboard, of which we use the official evaluation codebase.  [^2] https://github.com/huggingface/open_asr_leaderboard Throughout this section, we use the Whisper normalizer for English (Radford et al., 2023).
@@ 4.3.1 ASR Results
Long-form transcription. Table 2 reports WER values across 4 long-form datasets with sequences up to 2 hours: TED-LIUM (Hernandez et al., 2018), Meanwhile (Radford et al., 2023), Rev16 (Radford et al., 2023), and Earnings21 (Rio et al., 2021). We see that DSM-ASR outperforms streaming baselines and most of the non-streaming baselines, except for Parakeet.
@@ 4.4.1 TTS Results
DSM $54.8\pm 2.3$ $2048.3\pm 18.6$ $62.5\pm 2.0$ $2098.5\pm 21.1$
DSM (turn-by-turn) $53.1\pm 2.2$ $\mathbf{2092.6}\pm 18.3$ $61.9\pm 2.0$ $\mathbf{2179.1}\pm 22.4$
@@
Training on public data only. We provide results when using only public training data, and evaluating on Librispeech test clean (Panayotov et al., 2015) and on Seed test en (Anastassiou et al., 2024). We provide the results in the Appendix J, Tables 11 and 12, with comparisons to F5TTS (Chen et al., 2024b) and CosyVoice3-1.5B (RL)  (Du et al., 2025).
@@
DSM-ASR and DSM-TTS as a speech interface for LLMs. We combine DSM-ASR, DSM-TTS, and Gemma 3 (Gemma Team et al., 2025) into an LLM voice chat application with sub-second latency. We provide conversation samples at https://unmute.sh.
@@ Appendix I Latency, RTF, and throughput
F5-TTS (32 flow steps) 2.42 0.66 $\sim 400\,\text{ms}$ 7.1 7.1
DSM-TTS (ours, nq=16, b.s.=1) 750M 88K (EN) 1.95 0.67 139ms 4.4 4.4
DSM-TTS (ours, nq=32, b.s.=1) 900M 88K (EN) 1.68 0.71 172ms 2.7 2.7
DSM-TTS (ours, nq=32, b.s.=32) 900M 88K (EN) 1.68 0.71 351ms 2.3 74.2
Table 12: Results of DSM-TTS trained on public datasets, evaluated on Seed test en.
Model WER (%) $\downarrow$ SIM-o $\uparrow$
F5-TTS (Chen et al., 2024b) 1.83 0.71
Cosyvoice 3-1.5B (RL) (Du et al., 2025) 1.45 0.70
DSM-TTS (ours, nq=16) 1.58 0.70
DSM-TTS (ours, nq=32) 1.71 0.71
@@
We experiment with training DSM-TTS using only the publicly available datasets used for the short-form fine tuning of the ASR model, as described in Section A.

### v2

@@ Abstract
We introduce Delayed Streams Modeling (DSM), a flexible formulation for streaming, multimodal sequence-to-sequence learning. Sequence-to-sequence generation is often cast in an offline manner, where the model consumes the complete input sequence before generating the first output timestep. Alternatively, streaming sequence-to-sequence rely on learning a policy for choosing when to advance on the input stream, or write to the output stream. DSM instead models already time-aligned streams with a decoder-only language model. By moving the alignment to a pre-processing step, and introducing appropriate delays betweem streams, DSM provides streaming inference of arbitrary output sequences, from any input combination, making it applicable to many sequence-to-sequence problems. In particular, given a text and audio stream, automatic speech recognition (ASR) corresponds to the text stream being delayed, while the opposite gives a text-to-speech (TTS) model. We perform extensive experiments for these two major sequence-to-sequence tasks, showing that DSM provides state-of-the-art performance and latency while supporting arbitrary long sequences, being even competitive with offline baselines. Code, samples and demos are available at github.com/kyutai-labs/delayed-streams-modeling.
@@ 4.2 Training protocol
Finetuning (DSM-ASR). We then finetune the model on a collection of public datasets with ground-truth transcripts, described in Appendix A.1 and totaling 28k hours. This training stage lasts for 100k updates with batches of 128 examples, using 16 H100s. We augment training by using the codebook dropout (Défossez et al., 2024). We then adapt the model to long-form inputs, which most public datasets lack, by constructing a long-form mixture described in Appendix A.2. We run this stage for DSM-ASR for 25k updates with batch size 32, using 16 H100s.
@@ 4.3 Automatic Speech Recognition
Baselines. We benchmark DSM-ASR against leading models of the OpenASR Leaderboard, including Whisper (Radford et al., 2023), Canary-Flash (Zelasko et al., 2025), Phi-4 Multimodal Instruct (Abouelenin et al., 2025), Parakeet (Xu et al., 2023), Voxtral Mini (Liu et al., 2025) and as well as the closed-source ElevenLabs API 111elevenlabs.io.. Notably, all these models perform non-streaming ASR, as they require access to the full input sequence. We thus include a streaming variant of Whisper (Macháček et al., 2023) (with a delay of 2.5s) and SeamlessStreaming (Barrault et al., 2023). We also include a (block) streaming version of Parakeet-TDT-v2, evaluated via running the official scripts with a 2.5s delay that matches the delay used by DSM-ASR. [^2] We use chunk size of 10 frames (0.8s) and right context of 21 frames (1.7s). For long-form transcription, we add the Distil-Whisper (Gandhi et al., 2023) variant.
Transcription quality. We report micro-averaged Word Error Rate (WER), which avoids overemphasizing short sequences, and is the standard computation used in the OpenASR Leaderboard, of which we use the official evaluation codebase. Throughout this section, we use the Whisper normalizer for English (Radford et al., 2023).
@@ 4.3.1 ASR Results
Long-form transcription. Table 2 reports WER values across 4 long-form datasets with sequences up to 2 hours: TED-LIUM (Hernandez et al., 2018), Meanwhile (Radford et al., 2023), Rev16 (Radford et al., 2023), and Earnings21 (Rio et al., 2021). The non-streamed version of Parakeet runs out of memory on a 80Gb GPU even when we use single-example batches, so we exclude it from the analysis.
From Table 2, we see that DSM-ASR outperforms streaming and non-streaming baselines except for Parakeet, which it closely matches (7.8 vs 7.9).
@@ 4.4.1 TTS Results
DSM-TTS (ours) $54.8\pm 2.3$ $2048.3\pm 18.6$ $62.5\pm 2.0$ $2098.5\pm 21.1$
DSM-TTS (ours, turn-by-turn) $53.1\pm 2.2$ $\mathbf{2092.6}\pm 18.3$ $61.9\pm 2.0$ $\mathbf{2179.1}\pm 22.4$
@@
F5-TTS (32 flow steps) 2.42 0.66 $\sim 400\,\text{ms}$ 7.1 7.1
DSM-TTS (ours, nq=16, b.s.=1) 750M 88K (EN) 1.95 0.67 139ms 4.4 4.4
DSM-TTS (ours, nq=32, b.s.=1) 900M 88K (EN) 1.68 0.71 172ms 2.7 2.7
DSM-TTS (ours, nq=32, b.s.=32) 900M 88K (EN) 1.68 0.71 351ms 2.3 74.2
Table 8: Results of DSM-TTS trained on public datasets, evaluated on Seed test en. See Section J for details.
Model WER (%) $\downarrow$ SIM-o $\uparrow$
F5-TTS (Chen et al., 2024b) 1.83 0.71
Cosyvoice 3-1.5B (RL) (Du et al., 2025) 1.45 0.70
DSM-TTS (ours, nq=16) 1.58 0.70
DSM-TTS (ours, nq=32) 1.71 0.71
Training on public data only. For reproducibility, we provide results training from scratch DSM-TTS with the public training data described in Section J, totaling 88k hours of speech. We train a 300M parameters backbone model with either 16 or 32 codebooks. We evaluate on Librispeech test clean (Panayotov et al., 2015) and on Seed test en (Anastassiou et al., 2024). We provide the results in Tables 7 and 8, with comparisons to F5-TTS (Chen et al., 2024b) and CosyVoice3-1.5B (RL)  (Du et al., 2025). Refer to Appendix J for details.
@@
DSM-ASR and DSM-TTS as a speech interface for LLMs. We combine DSM-ASR, DSM-TTS, and Gemma 3 (Gemma Team et al., 2025) into an LLM voice chat application with sub-second latency. The application is available at https://unmute.sh.
@@ Appendix I Latency, RTF, and throughput
We experiment with training DSM-TTS using only the publicly available datasets.
@@
Dataset
We use a similar mixture as described in Section A.1, using the following datasets: AMI, EARNINGS22, GIGASpeech, SPGISpeech, TED-LIUM, VoxPopuli. Besides, we use LibriHeavy (Kang et al., 2023) and Emilia (He et al., 2024). For LibriHeavy, we use the original formatted text. We filter out data with a high word error rate compared with Whisper transcripts. This totals to 88k hours of speech.